# On the power of graph neural networks and the role of the activation function

### v5

@@ 3. Overview of the proofs
The proof of the first result is structured as follows. Since the trees are parameterized by $m$-tuples of integers $k_{1},\ldots,k_{m}$, the embedding of the root node computed by the GNN at any iteration is a function of these $m$ integers. Since the activations are piecewise polynomials, these embeddings of the root node are also piecewise multivariate polynomial functions of $k_{1},\ldots,k_{m}$. We further prove that the complexity of its polynomial pieces can be controlled uniformly in the sense of Definition 2: the number of generators of such polynomial is independent of $m$, but depends only on the number of iterations of the GNN and the underlying neural network (Lemma 2 and 3). Then, we show that there exists a large enough region of $\mathbb{R}^{m}$ on which this piecewise polynomial function is evaluated by the same polynomial. This region is large enough in the following sense: we prove that for a dimension $m$ that is sufficiently large, the region contains more integral vectors than the number of possible values a $q$-generated polynomial with degree at most $q$ can take on these vectors, even after identifying vectors up to permutations of the coordinates (Lemmas LABEL:lemmanbvaluesgeneratedp and 1). This implies that the polynomial piece will take the same value on two distinct integral vectors whose coordinates are not identical up to permutations. When translating this result back to the world of GNNs, this means that the two embeddings of the root nodes of the trees corresponding to these two vectors will coincide. To conclude a separation between bounded and unbounded GNNs, we justify that the unbounded ones can seperate these two vertices. This is based on the previous result (Theorem 2) stating that unbounded GNNs refine color refinement.
@@ 4. Collision with piecewise polynomial activations
Let $q$ be a positive integer. Then, there exists $m\in\mathbb{N}$ and two integral vectors $(k_{1},\cdots,k_{m})\in\mathbb{N}^{m}$ and $(k^{\prime}_{1},\cdots,k^{\prime}_{m})\in\mathbb{N}^{m}$ that are not equal up to a permutation such that for any sequence of symmetric piecewise polynomial functions $(f_{p}:\mathbb{R}^{p}\rightarrow\mathbb{R})_{p\in\mathbb{N}}$ satisfying $\mathsf{deg}(f_{p})\leq q$ for all $p\in\mathbb{N}$ (bounded degree condition on each polynomial piece for any $p$), $f_{m}(k_{1},\cdots,k_{m})=f_{m}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$.
@@
Let $m$ be any natural number such that $m>\max\{q^{2},2q\}$. For any natural number $M$, let $F_{M}$ be the box $\{(k_{1},\cdots,k_{m})\in\mathbb{Z}^{m}:\forall i\quad 1\leq k_{i}\leq M\}$. Let $\Omega_{M}:=\{\{\{x_{1},\cdots,x_{m}\}\}:\,(x_{1},\cdots,x_{m})\in F_{m}\}$; in other words, $\Omega_{M}$ is the set of multisets of size $m$ whose elements, when arranged as vectors, are in the box $F_{M}$ [^1] Another way to define $\Omega_{M}$ is as the orbits of the action of the symmetric group $\mathfrak{S}_{m}$ on $F_{M}$.. Consider
$\Phi:$ $\quad\Omega_{M}\longrightarrow\mathbb{Z}^{q}$
$\quad S\mapsto\,(\,s_{1}(S),s_{2}(S),\cdots,s_{q}(S)\,)$
which is well-defined because of the symmetry of the elementary symmetric polynomials $s_{1},\ldots,s_{q}$. Note that for any $i=1,\ldots,q$, $s_{i}$ is a sum of (${m\choose i}$ monomials whose maximum value is $M^{i}$ on $F_{M}$. Therefore, $\lvert\mathsf{Im}(\Phi)\rvert\leq(\sum_{i=1}^{q}\binom{m}{i}M^{i})^{q}\leq M^{q^{2}}(\binom{m}{q}q)^{q}$ where the last inequality follows from $\binom{m}{i}\leq\binom{m}{q}$ because $m>2q$.
On the other hand, $\lvert\mathsf{Dom}(\Phi)\rvert=\binom{M+m-1}{m}$ (number of multisets of size $m$ whose elements are taken from a set of size $M$). Now, let $f_{m}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ be the $m$-th function in the given sequence of symmetric piecewise polynomial functions where each polynomial piece has degree at most $q$. Let $r$ be the number of pieces of $f_{m}$. Then, there is a subset of $\mathsf{Dom}(\Phi)$ with at least $\frac{1}{r}\binom{M+m-1}{m}$ elements where $f_{m}$ is a symmetric polynomial $P$ of degree at most $q$. Lemma LABEL:lemma:symmultipol tells us that $P$ can be expressed as a polynomial of degree at most $q$ of the elementary symmetric polynomials $s_{1},\cdots,s_{q}$. Due to the pigeonhole principle, any such polynomial will be equal on two distinct multisets $\{\{k_{1},\cdots,k_{m}\}\}$ and $\{\{k^{\prime}_{1},\cdots,k^{\prime}_{m}\}\}$ in $\Omega_{M}$ as soon as:
@@
(1) $\underbrace{\frac{1}{r}\binom{M+m-1}{m}}_{\text{number of points}}\quad>\quad M^{q^{2}}(\binom{m}{q}q)^{q}\quad\geq\;\underbrace{\lvert\mathsf{Im}(\Phi)\rvert}_{\text{number of values $P$ can take at most}}$
@@
Such a value for $M$ can be found by noticing that $\binom{M+m-1}{m}$ is a polynomial of $M$ of degree $m$ whereas $M^{q^{2}}(\binom{m}{q}q)^{q}$ is a polynomial of $M$ of degree $q^{2}$. Since we chose $m$ to be greater than $q^{2}$, there exists $M\in\mathbb{N}$ such that Equation 1 is true. Hence there exist $k$ and $k^{\prime}$ whose coordinates are not equal up any permutation and such that $s_{i}(k_{1},\cdots,k_{m})=s_{i}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$ for any $i\in\{1,\cdots,m\}$. In turn $f_{m}(k_{1},\cdots,k_{m})=f_{m}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$.

### v6

@@ 3. Overview of the proofs
The proof of the first result is structured as follows. Since the trees are parameterized by $m$-tuples of integers $k_{1},\ldots,k_{m}$, the embedding of the root node computed by the GNN at any iteration is a function of these $m$ integers. Since the activations are piecewise polynomials, these embeddings of the root node are also piecewise multivariate polynomial functions of $k_{1},\ldots,k_{m}$. We further prove that the complexity of its polynomial pieces can be controlled uniformly in the sense of Definition 2: the number of generators of such polynomial is independent of $m$, but depends only on the number of iterations of the GNN and the underlying neural network (Lemma 3 and 4). Then, we show that there exists a large enough region of $\mathbb{R}^{m}$ on which this piecewise polynomial function is evaluated by the same polynomial. This region is large enough in the following sense: we prove that for a dimension $m$ that is sufficiently large, the region contains more integral vectors than the number of possible values a $q$-generated polynomial with degree at most $q$ can take on these vectors, even after identifying vectors up to permutations of the coordinates (Lemmas 1 and 2). This implies that the polynomial piece will take the same value on two distinct integral vectors whose coordinates are not identical up to permutations. When translating this result back to the world of GNNs, this means that the two embeddings of the root nodes of the trees corresponding to these two vectors will coincide. To conclude a separation between bounded and unbounded GNNs, we justify that the unbounded ones can seperate these two vertices. This is based on the previous result (Theorem 2) stating that unbounded GNNs refine color refinement.
@@ 4. Collision with piecewise polynomial activations
Let $m$, $q$ and $T$ be positive integers. For any natural number $M$, let $F_{M}$ be the box $\{(k_{1},\cdots,k_{m})\in\mathbb{Z}^{m}:\forall i\quad 1\leq k_{i}\leq M\}$ and let $\sigma_{1},\cdots,\sigma_{q}$ be $q$ reals. Let $P_{1},\cdots,P_{T}$ be $T$ polynomials of total degree less than $q$, and $q$-generated by $\sigma_{1},\cdots,\sigma_{q}$. Then the number of values taken by the function $x\mapsto(P_{1}(x),\cdots,P_{T}(x))$ on $F_{M}$ is at most
$\left(2\left(\max_{i\in[q],j\in[T],\alpha\in S}\lvert\lambda^{j}_{\alpha,i}\rvert\right)M^{q}\binom{m+q-1}{q}+1\right)^{qT}$
where $S$ is the union of the set of exponents of the $P_{j}$’s and the $\lambda^{j}_{\alpha,i}$ are the coefficients in the decomposition of their coefficients over the generators $\sigma_{1},\cdots,\sigma_{q}$.
@@
Note that if every $P_{j}$ is a polynomial with integer coefficients, i.e., each $P_{j}$ is generated by $1$, and has degree less than $q$, then the proof follows by considering the maximum and the minimum of the $P_{j}$’s over $F_{M}$, as each of them takes only integer values. To deal with the finitely generated case, we reduce it to the integer case. If $S$ is the union of the set of exponents of $P_{1},\ldots,P_{T}$, by assumption there exist integers $(\lambda^{j}_{\alpha,i})_{(\alpha,i)\in S\times\{1,\cdots,q\}}$ such that
$P_{j}=\sum_{\alpha\in S}\gamma_{\alpha,j}X^{\alpha}=$ $\sum_{\alpha\in S}\left(\sum_{i=1}^{q}\lambda^{j}_{\alpha,i}\sigma_{i}\right)X^{\alpha}$
$=$ $\sum_{i=1}^{q}\left(\sum_{\alpha\in S}\lambda^{j}_{\alpha,i}\sigma_{i}X^{\alpha}\right)$
$=$ $\sum_{i=1}^{q}\left(\underbrace{\sum_{\alpha\in S}\lambda^{j}_{\alpha,i}X^{\alpha}}_{:=f_{i,j}}\right)\sigma_{i}$
$f_{i,j}$ is a polynomial of degree less than $q$ with integer coefficients. As we wish to upper-bound the number of values of $\phi:x\mapsto(P_{1}(x),\cdots,P_{T}(x))$ on $F_{M}$, we consider
$\Phi:$ $\quad F_{M}\longrightarrow\mathbb{Z}^{qT}$
$\quad x\mapsto\,(\,f_{i,j}(x)\,)_{i\in[q],j\in[T]}$
The number of values of $\Phi$ on $F_{M}$ controls the number of values of $\phi$ on $F_{M}$ as each of the coordinates of $\phi$ is a linear combination of the coordinates of $\Phi$ with fixed linear coefficients $\sigma_{1},\ldots,\sigma_{q}$. We use the fact that each coordinate of $\Phi(x)$ is a multivariate polynomial with integer coefficients:
$\rvert\phi(F_{M})\rvert\leq\lvert\Phi(F_{M})\rvert$ $\leq\prod_{i\in[q],j\in[T]}(\lvert\max_{x\in F_{M}}f_{i,j}(x)-\min_{x\in F_{M}}f_{i,j}(x)\rvert+1)$
$\leq\prod_{i\in[q],j\in[T]}(2\max_{x\in F_{M}}\lvert f_{i,j}(x)\rvert+1)$
$\leq\left(2\left(\max_{i\in[q],j\in[T],\alpha\in S}\lvert\lambda^{j}_{\alpha,i}\rvert\right)M^{q}\binom{m+q-1}{q}+1\right)^{qT}$
where the last inequality follows from the fact that $\binom{m+q-1}{q}$ is an upper-bound on the number of monomials of $f_{i,j}$. ∎
Lemma 2.
Let $q$ be a positive integer, let $I$ be a finite subset of $\mathbb{N}$, and let $(f_{t,m}:\mathbb{R}^{m}\rightarrow\mathbb{R})_{(t,m)\in I\times\mathbb{N}}$ be a double sequence of piecewise polynomial functions satisfying:
  • i)
    $\mathsf{deg}(f_{t,m})\leq q$ for all $(t,m)\in I\times\mathbb{N}$ (bounded degree condition).
  • ii)
    for any $m\in\mathbb{N}$ and $t\in I$, $f_{t,m}$ is $q$-generated.
Then, there exists $m\in\mathbb{N}$ and two integral vectors $(k_{1},\cdots,k_{m})\in\mathbb{N}^{m}$ and $(k^{\prime}_{1},\cdots,k^{\prime}_{m})\in\mathbb{N}^{m}$ that are not equal up to a permutation such that for any $t\in I$, $f_{t,m}(k_{1},\cdots,k_{m})=f_{t,m}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$.
Proof.
Let $q$ be a positive integer, $I$ be a finite subset of $\mathbb{N}$ and $(f_{t,m}:\mathbb{R}^{m}\rightarrow\mathbb{R})_{(t,m)\in I\times\mathbb{N}}$ be a double sequence of piecewise polynomial satisfying the assumptions of the lemma. Let $m$ be any natural number such that $m>q^{2}\lvert I\rvert$. For any positive integer $M$, let $F_{M}$ be the box $\{(k_{1},\cdots,k_{m})\in\mathbb{Z}^{m}:\forall i\quad 1\leq k_{i}\leq M\}$.
There are $\binom{M+m-1}{m}$ multisets of size $m$ that can be formed from $\{1,\cdots,M\}$. One can find that many elements of $F_{M}$ that are not equal up to a permutation. Let $r$ be an upperbound on the number of pieces of $f_{t,m}$ for every $t$, i.e. $r:=\max\{\text{number of pieces of }f_{t,m}:t\in I\}$: note that $r$ exists since $I$ is finite, but may depend on $m$. Then, there exists a subset of $F_{M}$ with at least $\frac{1}{r^{\lvert I\rvert}}\binom{M+m-1}{m}$ integral vectors that are not equal up to a permutation where for every $t$, $f_{t,m}$ is a polynomial $P_{t,m}$ of degree at most $q$, and $P_{t,m}$ is $q$-generated. This is true as the collection of piecewise polynomial functions $(f_{i,m})_{i\in I}$ divide $F_{M}$ in at most $r^{\lvert I\rvert}$ regions, where in each region, for every $i\in I$, $f_{i,m}$ is polynomial.
Due to Lemma 1, the pigeonhole principle gives us that all polynomials will be equal on some vectors $(k_{1},\cdots,k_{m})$ and $(k^{\prime}_{1},\cdots,k^{\prime}_{m})$ of $F_{M}$ not equal up to a permutation as soon as:
@@
(1) $\underbrace{\frac{1}{r^{\lvert I\rvert}}\binom{M+m-1}{m}}_{\begin{subarray}{c}\text{number of multisets formed}\\ \text{ from a region of $F_{M}$ where }\\ \text{ each $f_{t,m}$ is polynomial }\end{subarray}}\quad>\underbrace{\left(2\max_{i\in[q],t\in I}(\max_{\alpha\in S_{m,t}}\lvert\lambda^{m,t}_{\alpha,i}\rvert)M^{q}\binom{m+q-1}{q}+1\right)^{q\lvert I\rvert}}_{\begin{subarray}{c}\text{number of values $(P_{1,m},P_{2,m},\cdots,P_{\lvert I\rvert,m})$}\\ \text{ can take at most on $F_{M}$}\end{subarray}}$
where the $\lambda^{m,t}_{\alpha,i}$ are integers in the decomposition of the polynomials $f_{t,m}$ over generators $\sigma^{m}_{1},\cdots,\sigma^{m}_{q}\in\mathbb{R}$ for the monomial $X^{\alpha}$.
@@
Such a value of $M$ can be found by noticing that $\binom{M+m-1}{m}$ is a polynomial of $M$ of degree $m$ whereas the right hand side is a polynomial of $M$ of degree $q^{2}\lvert I\rvert$. Since we chose $m$ to be greater than $q^{2}\lvert I\rvert$, there exists $M\in\mathbb{N}$ such that Equation (1) holds. Hence there exist $k$ and $k^{\prime}$ whose coordinates are not equal up any permutation and such that $f_{i,m}(k_{1},\cdots,k_{m})=f_{i,m}(k^{\prime}_{1},\cdots,k^{\prime}_{m})$ for any $i\in I$. ∎